OpenStudy (anonymous):

The perimeter of a triangle is 45 centimeters. The longest side is 5 centimeter(s) less than the sum of the other two sides. Twice the shortest side is 4 centimeters less than the longest side. Find the lengths of each side of the triangle. The length from the shortest to longest are_, _, and _ centimeters?

OpenStudy (anonymous):

The information you're given allows you to set up three equations in three unknowns. Name the sides as: l = longest s = shortest a = remaining side The perimeter is 45, which is the sum of the lengths of the sides, so\[l+s+a=45\]is one equation. You're told that the longest side is 5 LESS than the SUM of the other two sides, so\[l=a+s-5\]and finally, TWICE the shortest side equals 4 LESS than the longest, namely,\[2s=l-4\]

OpenStudy (anonymous):

You have three linear equations in three unknowns, so you can solve the system exactly.

OpenStudy (anonymous):

Lol. Ok, so what;s the answers though?

OpenStudy (anonymous):

i like the way you explain it though...

OpenStudy (anonymous):

Well, you have to do some substituting. Take any one of you equations to begin with. Use the second to plug into the third to eliminate for l. So,\[2s=(a+s-5)-4=a+s-9 \rightarrow a=s+9\]

OpenStudy (anonymous):

Let me sort it.

OpenStudy (anonymous):

.......

OpenStudy (anonymous):

It's fast to do on paper but a pain to do online.

OpenStudy (anonymous):

tell me about it...

OpenStudy (anonymous):

If you sub. the equation for l (second one) into the first, you should get\[(a+s-5)+s+a=45 \rightarrow s+a=20\]

OpenStudy (anonymous):

If you sub the same second equation for l into the third, you get\[2s=(a+s-5)-4 \rightarrow a-s=9\] (forget the equation I solved at the very beginning, I stuffed the sign).

OpenStudy (anonymous):

You now have \[a+s=20\]and\[a-s=9\]

OpenStudy (anonymous):

If a=9+s from the second equation here, then putting it into the first, you get\[(9+s)+s=20 \rightarrow 2s=11 \rightarrow s=\frac{11}{2}\]

OpenStudy (anonymous):

lol..this is a monster!

OpenStudy (anonymous):

Which means, from any of the two equations involving a and s above, for a,\[a=9+s=9+\frac{11}{2}=\frac{29}{2}\]

OpenStudy (anonymous):

Now from your perimeter equation, you can solve for l:\[l=45-a-s=45-\frac{29}{2}-\frac{11}{2}=25\]

OpenStudy (anonymous):

You're done.

OpenStudy (anonymous):

Well, I am...

OpenStudy (anonymous):

so it's how many centimeters each?

OpenStudy (anonymous):

Just look through - l = 25 cm s = 5.5 cm a = 14.5 cm

OpenStudy (anonymous):

crap

OpenStudy (anonymous):

crap what?

OpenStudy (anonymous):

The correct answers were 8,17, and 20! :(

OpenStudy (anonymous):

Now it reformulates a new, but similar problem for me to solve.

OpenStudy (anonymous):

Now it's......

OpenStudy (anonymous):

The perimeter of a triangle is 45 centimeters. The longest side is 5 centimeter(s) less than the sum of the other two sides. Twice the shortest side is 4 centimeters less than the longest side. Find the lengths of each side of the triangle. The length from the shortest to longest are_, _, and _ centimeters?

OpenStudy (anonymous):

Yes a mistake was made. You need to check everything once it's solved but putting it back into the equations.

OpenStudy (anonymous):

its the same one! never had that happen before.

OpenStudy (anonymous):

so what would be the "mistake free" answer? :)

OpenStudy (anonymous):

Well, if it's the same question, you already have the answers - you just posted them!

OpenStudy (anonymous):

LOL......DUH!

OpenStudy (anonymous):

I didn't even think of that.

OpenStudy (anonymous):

Sigh......long day.

OpenStudy (anonymous):

Want a shot at another?

OpenStudy (anonymous):

Depends, if you give me back my fan point ;)

OpenStudy (anonymous):

Fair enough!

OpenStudy (anonymous):

Do you have the question?

OpenStudy (anonymous):

Yeah, it's a beast.... one sec.....

OpenStudy (anonymous):

In a certain year out of pocket expenses for an elderly person in poor health were 3353 more than for an elderly person in good health. Combined out of pocket expenses for one person in poor health and one person in good health was 6677 find the our of pocket expenses for each type of person. Out of pocket expenses for a person in poor health was.... Out of pocket expenses for a person in good health was....

OpenStudy (anonymous):

phew!

OpenStudy (anonymous):

Do you just want the answer?

OpenStudy (anonymous):

Did you take my point off again? lol

OpenStudy (anonymous):

lol...me? Nooooooo! Why would you ask that?! ;)

OpenStudy (anonymous):

poor health person = 5015 good health person = 1662

OpenStudy (anonymous):

Yes.....I was holding it hostage until you returned. It's back.

OpenStudy (anonymous):

Ok......now enlighten me how you got it so darn fast! I have a friend on here that is taking the same class and I want to help her too.

OpenStudy (anonymous):

She is Kelley29 I think..

OpenStudy (anonymous):

poor health out-of-costs = p good health out-of-costs = g Information says p = 3353 + g p + g = 6677 Solve the two. First is p=3353 + g so sub. into second to get rid of p: (3353 + g) = 6677

OpenStudy (anonymous):

.............

OpenStudy (anonymous):

g=6677-3353=1662

OpenStudy (anonymous):

p=3353+ g = 3353 + 1662 = 5015

OpenStudy (anonymous):

Ok what her numbers are poor health=3334, good health=6596

OpenStudy (anonymous):

she just gave me that, and I have no idea what that formula is you just gave me.....unfortunately. I will tell her to give you a POINT!

OpenStudy (anonymous):

you're a star

OpenStudy (anonymous):

Whats tha t mean?

OpenStudy (anonymous):

It just means I appreciate your offer to get your friend to give me a point.

OpenStudy (anonymous):

Lol, i'ts the least I can do!

OpenStudy (anonymous):

SO what is it?

OpenStudy (anonymous):

what is what?

OpenStudy (anonymous):

her answer! lol

OpenStudy (anonymous):

U need more info?

OpenStudy (anonymous):

I don't have her question - what's her login name?

OpenStudy (anonymous):

Just give me a second, I'm just helping someone else too.

OpenStudy (anonymous):

Same questions! Hers is Kelley29

OpenStudy (anonymous):

Ok n/m some dude just got it for her.

OpenStudy (anonymous):

But...we have more. At least a coupla more.

OpenStudy (anonymous):

Kelly29 is being helped by blexting.

OpenStudy (anonymous):

He/she does a good job.

OpenStudy (anonymous):

Ok, well I have more...

OpenStudy (anonymous):

Translate to a system of equations, but do not solve. An non toxic floor wax can be made from lemon juice and food grade linseed oil. The amount of oil should be twice the amount of lemon juice. How much of each ingredient is needed to make 42 ounces of floor wax? Let x represent the number of ounces of lemon juice and y the number of ounces of linseed oil. Complete the system of equations. y= x+y=

OpenStudy (anonymous):

It's absolutely ridiculous!

OpenStudy (anonymous):

...........

OpenStudy (anonymous):

y=2x x+y =42 ------------------------------------------------------ You're told you need 2 times as much x as y (2 times as much lemon juice as linseed oil). You also know that the total weight, x + y, should be 42 oz.

OpenStudy (anonymous):

ANd AGAIN........you got it right!

OpenStudy (anonymous):

last one here for the night.....i promise.........

OpenStudy (anonymous):

If a plane can travel 480 miles per hour with the wind, and 380 miles per hour against the wind, find the speed of the plan without a wind and the speed of the wind. The speed of the plan in still air is _miles per hour.

OpenStudy (anonymous):

It's 430 miles per hour

OpenStudy (anonymous):

Ok........the speed of the WIND is?

OpenStudy (anonymous):

50, sorry, forgot

OpenStudy (anonymous):

50mph?

OpenStudy (anonymous):

Ok, can you PLEASE help my friend now.....you rock, and these other guys suck apparently! I will tell her to rock your points! :,)

OpenStudy (anonymous):

Thank you very much by the way......I really do appreciate your help.

OpenStudy (anonymous):

Thanks for the praise. I'll see what I can do. I promised someone else I'd help him and he's been waiting. If kelley29 is still around when I'm done, I'll help.

OpenStudy (anonymous):

I understand. You're the best! I hope to see you on here for the next 4 weeks or so.